Write the slope intercept equation for a line passing through the given point that is parallel to the given line.
(3, 5), 𝑦=2/7𝑥+1

Found 2 solutions by josgarithmetic, lwsshak3:
Answer by josgarithmetic(39632)

Already answered here: http://www.algebra.com/algebra/college/linear/Linear_Algebra.faq.question.1001691.html

Answer by lwsshak3(11628)

Write the slope intercept equation for a line passing through the given point that is parallel to the given line.
(3, 5), 𝑦=2/7𝑥+1
Equation of a line in slope-intercept form:y=mx+b, m=slope, b=y-intercept
y=(2/7)x+1
slope,m=2/7
Parallel lines have the same slope
y=(2/7)x+b
solve for b using given coordinates(3,5) on the line
b=y-(2/7)x=5-(2/7)*3
5-6/7=35/7-6/7=29/7
equation of line: y=(2/7)x+29/7